# Evaluating machine learning models for predicting pesticides toxicity to honey bees

### v2

@@ 2.3.5 Pretrained neural networks
Those models can be finetuned, using the pretraining model as a weight initialization 69. However, this approach assumes that there is enough data available to avoid overfitting the model, which is often not the case for highly specific domains in chemistry. Instead, one can use the molecule embeddings from the pretrained model, like a molecular fingerprint, and input it to any conventional ML model. This approach is equivalent to freezing most model weights and training only the classification head. It is also known as frozen embeddings 69, and has shown results competitive with finetuning models 70, while being faster and less prone to overfitting.
@@
Molecule Attention Transformer 65 and its later extension, the Relative Molecule Self-attention Transformer (R-MAT) 71, incorporate many inductive biases specific to molecular chemistry into the GNN graph transformer and their pre-training process. They are based on self-attention between atoms, which is a weighted sum of three components, incorporating molecule-specific inductive biases. In particular, both utilize key-value attention between atom embeddings and pairwise shortest paths between atoms, incorporating both local and global molecule context. Both follow the encoder-only transformer architecture (similar to BERT 72), interleaving layers of attention, normalization, and MLPs. The learned atom embeddings are combined with the average readout in MAT and with the attention-based pooling in R-MAT. The pretraining procedure for MAT is based only on masked node reconstruction 73, which is a graph-based equivalent of masked language modeling, inspired by BERT. R-MAT uses more challenging masked subgraph reconstruction (similar to GROVER 74), and multitask regression at the molecule level (inspired by MolBERT 70).
@@
GROVER 74 is a GNN based on a graph transformer architecture. It creates embeddings for both atoms and bonds in parallel for increased expressivity, and includes skip connections and randomized message passing to better learn long-range relations in molecules. It proposed using two pretraining tasks: masked subgraph reconstruction, and classification of functional groups existence. This incorporates strong subgraph-level knowledge.
@@
A notable family of novel molecular models are transformers based solely on SMILES, as exemplified by ChemBERTa 75, 76. It follows the decoder-only architecture and can be pretrained only on masked language modeling (MLM) on SMILES. Alternatively, it can be pretrained on multitask regression (MTR), predicting many physicochemical properties of molecules computed with RDKit. BERT-like models have relatively simple architectures but result in highly performant embeddings.
@@
Lastly, a hybrid model between graph-based and SMILES-based models is Mol2Vec 77. It treats the molecule as a sequence of words, based on ECFP invariants 23, computed from the molecular graph. Then it vectorizes them like words, using a pretrained Word2Vec model. Those embeddings are then added to obtain the embedding for the whole molecule.
@@ 2.3.6 Explainability
We focus on a local, model-agnostic interpretability methods 78, and specifically counterfactual explanations. Those methods interpret a prediction made for a single molecule with an easy visualization of the results. They are independent of the model used, which is very useful as we explore a wide variety of approaches to molecular property prediction. For binary classification, as in ApisTox, counterfactual explanations are very useful, answering the following question: “What is the minimal change required for the input molecule so that the model changes the prediction to the opposite class?”. Algorithms of this type can suggest substituents, functional groups, or fragment removal from the input molecule, so that the model changes the prediction, e.g. from non-toxic to toxic.
Molecular Model Agnostic Counterfactual Explanations (MMACE) 79 has been designed as a fast and easy-to-use method in this group. Importantly, it does not require dataset-specific training, instead relying on directly modifying the molecule being explained using the STONED 80 generative model. In this way, it explored the chemical space around it, using only model predictions to determine samples from the opposite class. Tanimoto similarity on ECFP4 fingerprints is used to select counterfactuals most similar to the original molecule, that is, with the smallest change.
@@ 2.4.1 Models implementation
We implement molecular fingerprints using the scikit-fingerprints library 81, using all 2D fingerprints available in the library version 1.12.0. We omit 3D conformer-based fingerprints, since they gave very subpar results in initial experiments. For all fingerprints, we use Random Forest classifier from scikit-learn 82, with 100 trees (default value) and entropy as split criterion. Entropy worked slightly better than the default Gini entropy in our initial experiments, in agreement with previous works 33, 34, 83. For model regularization, we set the minimal number of samples to split.
Atom counts baseline 32 is implemented using RDKit, with Random Forest like for fingerprints. For LTP 33 and MOLTOP 34 baselines, we use the original code and classifier settings. Both use Random Forest, but with default settings tuned over a large number of datasets.
Graph kernels are implemented using GraKeL 84, with the kernel SVM classifier from scikit-learn. After initial experiments, we decided to use the following kernels: shortest paths, propagation, WL, WL-OA. Molecular graphs are attributed with atomic numbers and bond types (single, double, triple, aromatic, other). We tune kernel-specific hyperparameters (see supplementary information for details) and SVM regularization $C$.
GNNs are implemented in PyTorch 85 and PyTorch Geometric 86. We implement GCN, GraphSAGE, GIN, GAT, and AttentiveFP. Following many architectures in a popular OGB benchmark 87, we extract a standard set of features for atoms and bonds, and use the embedding layer to obtain their initial vectors. Atom features are based on invariants used in ECFP, e.g. atomic number, degree, formal charge. However, we note that only GIN, GAT, and AttentiveFP make use of bond features. We also use Jumping Knowledge (concatenation variant) 62, LayerNorm 63, 64. For readout, we use sum pooling, proven to be the most expressive 59. We make an exception for AttentiveFP, which uses its original GRU-based pooling instead. We use the same number of channels in each layer and treat the number of layers and channels as hyperparameters, while also tuning the learning rate and dropout. Networks are trained for 100 epochs, without early stopping.
For pretrained models, we use MAT, R-MAT, GROVER, ChemBERTa, and Mol2Vec. All are used as embedding models, similar to fingerprints. We do not perform fine-tuning, since during initial experiments it constantly overfitted and gave worse result than the frozen embeddings approach. Initial experiments indicated that logistic regression performs considerably better than Random Forest as a classifier in this case, probably because pretraining used such linear classifiers and the feature space is better aligned for them. Therefore, we use logistic regression as classifier and tune the regularization strength $C$.
@@ 2.4.2 Evaluation
For a realistic and challenging train-test split, we use two splits provided with ApisTox: MaxMin split and time split approximation. The test set in MaxMin split is created by selecting maximally diverse molecules, i.e. with the highest sum of pairwise Tanimoto distances between their ECFP4 fingerprints. This way, we require the toxicity classifier to perform well across the whole chemical space of the dataset. Since the maximum diversity picking problem is NP-hard 88, MaxMin heuristic from RDKit is used here 16. Time split is approximated by PubChem literature data, with the test set consisting of the newest agrochemicals. This requires out-of-distribution generalization capabilities, since new pesticides are typically structurally novel. For both splits, the test set is 20% of the data.
To validate the performance of the model, we measure multiple metrics, appropriate for imbalanced datasets: AUROC, precision, recall, and Matthews correlation coefficient (MCC) 89, 90, 91. Comparison of models based on multiple metrics is important, as they highlight potential differences in their performance. In particular, MCC is reported to better reflect model performance compared to AUROC 89, and precision and recall focus on false positives and false negatives, respectively.
@@
We include the hyperparameter ranges and other details in the supplementary information.
@@ 2.5 Molecular dataset analysis
RDKit 92 was used to calculate molecular descriptors and most chemoinformatic analyzes. scikit-fingerprints 81 was used to calculate molecular fingerprints and diversity calculations.
#Circles algorithm was implemented using the description from the original publication 27, as well as the official implementation. As the exact computation has exponential complexity, we use the fast sequential approximation proposed in the original implementation. We use the same parameters as the original publication, i.e. threshold 0.75 of Tanimoto distance and ECFP4 fingerprint of length 1024 as the feature space.
@@ 2.6 Explainability
We use the original MMACE implementation from the ExMol package 79. We generate 5000 samples and do not use drug-like filtering (used in the original MMACE paper), in order to properly explore the pesticides chemical space. To avoid data leakage, we explain only the model predictions on molecules from the test set.
@@ 3 Data and software availability
We release our code publicly at https://github.com/j-adamczyk/ApisTox_bee_toxicity_ML_prediction. The ApisTox dataset, along with MaxMin and time splits, is available on Zenodo 93 and GitHub 94. We also include it in our GitHub repository with code for reproducing experimental results.
@@ 4 Results and discussion
The main results and quality metrics are reported in Table 1 (MaxMin split) and Table 2 (time split). For brevity, we include the top 5 fingerprints (with the highest MCC scores), with results for the rest in the supplementary information. The best metric value in each group is marked in bold (two values in case of ties).
@@
Molecular fingerprints are, overall, the most performant and robust solution. Interestingly, quite different fingerprints are the best for both splits. For MaxMin, the best fingerprint is Avalon 95, with 3 substructural fingerprints (Laggner 20, MACCS 21, PubChem 24) also among the top 5. However, for time split, all best fingerprints are hashed, e.g. ECFP 23, Layered 92, RDKit 92. This confirms that flexible, more data-driven hashed fingerprints are better suited for out-of-distribution generalization. In contrast, in-domain predictions can be made accurately by extracting common functional groups and other substructural patterns. Since a very large number of fingerprint algorithms are available, understanding such implications can be useful for practical applications.
The performance of GNNs trained from scratch is very underwhelming in both cases. All are unable to outperform baselines, even the simplest atom count in the MaxMin split. This is despite significant computational resources and extensive hyperparameter tuning, even for AttentiveFP, designed specifically for molecular property prediction. However, GNNs are known to require larger amounts of data to train effectively, and by its nature, agrochemistry requires low-data ML. Furthermore, it frequently involves salts, mixtures, and other molecules with disconnected components in their molecular graphs 14, 96, 97. This is problematic for message-passing models, since they can obtain this information only in the last readout layer, limiting the effectiveness of node embedding learning from their neighborhoods.
@@ 4.1.3 Share of fragmented molecules
Fragmented molecules are defined as those in which two or more independent fragments are present in a single SMILES record. Examples include salts or mixtures, frequently found in agrochemical QSAR 98, 99. Such compounds can be challenging for some methods, especially those based on graph neural networks (GNNs) or path-based fingerprints like Atom Pair. In such cases, the individual fragments are treated as separate graphs, which prevents the message passing between them. This can impact the performance of the models.
@@ 4.1.4 Unique functional groups
The functional groups defined in the Laggner fingerprint 20 were utilized. It is a collection of 307 fragments, also including those typically indicating toxicity or generally less common in typical drug-like collections.
@@ 4.1.5 Chemical diversity
Furthermore, all analyzed datasets appear to be relatively distant from each other (average Tanimoto similarity $\leq 0.56$), suggesting that they represent different areas of chemical space. However, it should be noted that the results obtained from this analysis show little correlation with the difficulty of their respective classification tasks and the quality of the resulting ML models 34.
@@ 4.2 Interpretability - counterfactual explanations
Overall, the most prevalent motifs influencing model predictions are connected to polarity, ionization states, charge distribution, and stability (e.g. ester bonds). In fact, these are some of the key factors that organic chemists and environmental scientists consider when designing molecules. The polarity and ionization states influence solubility and control the mobility of the compound in the environment, affecting aspects like runoff, leaching into groundwater, and bioavailability. More polar molecules tend to have lower bioaccumulation potential, as they interact less strongly with the lipid membranes of organisms 100. Charge characteristics can dictate how a molecule adsorbs in soils or sediments, affecting its overall persistence and exposure risks 101. Finally, incorporating chemically labile groups like ester bonds is a common strategy to design molecules that break down into less harmful components after they have performed their intended function. This is part of a “benign-by-design” approach to reduce long-term environmental impact 102.
@@ 4.3 Interpretability - uncertain molecules
In addition to the main dataset and its splits, ApisTox provides a smaller dataset of molecules rejected from the main dataset due to disagreements between experimental measurements in the ECOTOX database. For the main dataset, only molecules with perfectly agreeing $LD_{50}$ values were included, i.e., all under or over $11$ $\mu\text{g / bee}$ (US EPA guidelines threshold 103). Others are provided as a file with uncertain molecules. Performing predictions on those compounds provides additional evidence, in addition to the literature and experiments, about their toxicity. This is also a very practical application of proposed models, as conflicting measurements can easily arise in ecotoxicology research 104, 105, e.g. due to variations in experimental procedures or outliers in results.
@@
Indoxacarb, also known commercially as Steward, Avaunt, Advion, and Arilon, is an oxadiazine insecticide, used in many household pesticides, and is also approved for use in the US and UK. However, the EU withdrew its approval for the use of Indoxacarb as a plant protection insecticide at the end of 2021, due to its environmental impact and high toxicity. In particular, “a high risk to bees was identified for the representative use in maize, sweet corn and lettuce for seed production” 106. Thus, this is a rather nuanced situation with regulatory disagreements, and prediction of machine learning further corroborates the high toxicity of Indoxacarb toward honey bees.
Phosalone has values ranging from $9.9$ to $17.4$ in the data. The model predicts toxicity (positive class), which is in agreement with the EU decision from 2006 107.

### v3

@@ 2.3.5 Pretrained neural networks
Those models can be finetuned, using the pretraining model as a weight initialization [69]. However, this approach assumes that there is enough data available to avoid overfitting the model, which is often not the case for highly specific domains in chemistry. Instead, one can use the molecule embeddings from the pretrained model, like a molecular fingerprint, and input it to any conventional ML model. This approach is equivalent to freezing most model weights and training only the classification head. It is also known as frozen embeddings [69], and has shown results competitive with finetuning models [70], while being faster and less prone to overfitting.
@@
Molecule Attention Transformer [65] and its later extension, the Relative Molecule Self-attention Transformer (R-MAT) [71], incorporate many inductive biases specific to molecular chemistry into the GNN graph transformer and their pre-training process. They are based on self-attention between atoms, which is a weighted sum of three components, incorporating molecule-specific inductive biases. In particular, both utilize key-value attention between atom embeddings and pairwise shortest paths between atoms, incorporating both local and global molecule context. Both follow the encoder-only transformer architecture (similar to BERT [72]), interleaving layers of attention, normalization, and MLPs. The learned atom embeddings are combined with the average readout in MAT and with the attention-based pooling in R-MAT. The pretraining procedure for MAT is based only on masked node reconstruction [73], which is a graph-based equivalent of masked language modeling, inspired by BERT. R-MAT uses more challenging masked subgraph reconstruction (similar to GROVER [74]), and multitask regression at the molecule level (inspired by MolBERT [70]).
@@
GROVER [74] is a GNN based on a graph transformer architecture. It creates embeddings for both atoms and bonds in parallel for increased expressivity, and includes skip connections and randomized message passing to better learn long-range relations in molecules. It proposed using two pretraining tasks: masked subgraph reconstruction, and classification of functional groups existence. This incorporates strong subgraph-level knowledge.
@@
A notable family of novel molecular models are transformers based solely on SMILES, as exemplified by ChemBERTa [75, 76]. It follows the decoder-only architecture and can be pretrained only on masked language modeling (MLM) on SMILES. Alternatively, it can be pretrained on multitask regression (MTR), predicting many physicochemical properties of molecules computed with RDKit. BERT-like models have relatively simple architectures but result in highly performant embeddings.
@@
Lastly, a hybrid model between graph-based and SMILES-based models is Mol2Vec [77]. It treats the molecule as a sequence of words, based on ECFP invariants [23], computed from the molecular graph. Then it vectorizes them like words, using a pretrained Word2Vec model. Those embeddings are then added to obtain the embedding for the whole molecule.
@@ 2.3.6 Explainability
We focus on a local, model-agnostic interpretability methods [78], and specifically counterfactual explanations. Those methods interpret a prediction made for a single molecule with an easy visualization of the results. They are independent of the model used, which is very useful as we explore a wide variety of approaches to molecular property prediction. For binary classification, as in ApisTox, counterfactual explanations are very useful, answering the following question: “What is the minimal change required for the input molecule so that the model changes the prediction to the opposite class?”. Algorithms of this type can suggest substituents, functional groups, or fragment removal from the input molecule, so that the model changes the prediction, e.g. from non-toxic to toxic.
Molecular Model Agnostic Counterfactual Explanations (MMACE) [79] has been designed as a fast and easy-to-use method in this group. Importantly, it does not require dataset-specific training, instead relying on directly modifying the molecule being explained using the STONED [80] generative model. In this way, it explored the chemical space around it, using only model predictions to determine samples from the opposite class. Tanimoto similarity on ECFP4 fingerprints is used to select counterfactuals most similar to the original molecule, that is, with the smallest change.
@@ 2.4.1 Models implementation
We implement molecular fingerprints using the scikit-fingerprints library [81], using all 2D fingerprints available in the library version 1.12.0. We omit 3D conformer-based fingerprints, since they gave very subpar results in initial experiments. For all fingerprints, we use Random Forest classifier from scikit-learn [82], with 100 trees (default value) and entropy as split criterion. Entropy worked slightly better than the default Gini entropy in our initial experiments, in agreement with previous works [33, 34, 83]. For model regularization, we set the minimal number of samples to split.
Atom counts baseline [32] is implemented using RDKit, with Random Forest like for fingerprints. For LTP [33] and MOLTOP [34] baselines, we use the original code and classifier settings. Both use Random Forest, but with default settings tuned over a large number of datasets.
Graph kernels are implemented using GraKeL [84], with the kernel SVM classifier from scikit-learn. After initial experiments, we decided to use the following kernels: shortest paths, propagation, WL, WL-OA. Molecular graphs are attributed with atomic numbers and bond types (single, double, triple, aromatic, other). We tune kernel-specific hyperparameters (see A for details) and SVM regularization $C$.
GNNs are implemented in PyTorch [85] and PyTorch Geometric [86]. We implement GCN, GraphSAGE, GIN, GAT, and AttentiveFP. Following many architectures in a popular OGB benchmark [87], we extract a standard set of features for atoms and bonds, and use the embedding layer to obtain their initial vectors. Atom features are based on invariants used in ECFP, e.g. atomic number, degree, formal charge. However, we note that only GIN, GAT, and AttentiveFP make use of bond features. We also use Jumping Knowledge (concatenation variant) [62], LayerNorm [63, 64]. For readout, we use sum pooling, proven to be the most expressive [59]. We make an exception for AttentiveFP, which uses its original GRU-based pooling instead. We use the same number of channels in each layer and treat the number of layers and channels as hyperparameters, while also tuning the learning rate and dropout. Networks are trained for 100 epochs, without early stopping.
For pretrained models, we use MAT, R-MAT, GROVER, ChemBERTa, and Mol2Vec. All are used as embedding models, similar to fingerprints. We do not perform fine-tuning, since during initial experiments it constantly overfitted and gave worse result than the frozen embeddings approach. Initial experiments indicated that logistic regression performs considerably better than Random Forest as a classifier in this case, probably because pretraining used such linear classifiers and the feature space is better aligned for them. Therefore, we use logistic regression as classifier and tune the regularization strength $C$. For further details on those models, see C.
@@ 2.4.2 Evaluation
For a realistic and challenging train-test split, we use two splits provided with ApisTox: MaxMin split and time split approximation. The test set in MaxMin split is created by selecting maximally diverse molecules, i.e. with the highest sum of pairwise Tanimoto distances between their ECFP4 fingerprints. This way, we require the toxicity classifier to perform well across the whole chemical space of the dataset. Since the maximum diversity picking problem is NP-hard [88], MaxMin heuristic from RDKit is used here [16]. Time split is approximated by PubChem literature data, with the test set consisting of the newest agrochemicals. This requires out-of-distribution generalization capabilities, since new pesticides are typically structurally novel. For both splits, the test set is 20% of the data.
To validate the performance of the model, we measure multiple metrics, appropriate for imbalanced datasets: AUROC, precision, recall, and Matthews correlation coefficient (MCC) [89, 90, 91]. Comparison of models based on multiple metrics is important, as they highlight potential differences in their performance. In particular, MCC is reported to better reflect model performance compared to AUROC [89], and precision and recall focus on false positives and false negatives, respectively.
@@
We include the hyperparameter ranges and other details in the A.
@@ 2.5 Molecular dataset analysis
RDKit [92] was used to calculate molecular descriptors and most chemoinformatic analyzes. scikit-fingerprints [81] was used to calculate molecular fingerprints and diversity calculations.
#Circles algorithm was implemented using the description from the original publication [27], as well as the official implementation. As the exact computation has exponential complexity, we use the fast sequential approximation proposed in the original implementation. We use the same parameters as the original publication, i.e. threshold 0.75 of Tanimoto distance and ECFP4 fingerprint of length 1024 as the feature space.
@@ 2.6 Explainability
We use the original MMACE implementation from the ExMol package [79]. We generate 5000 samples and do not use drug-like filtering (used in the original MMACE paper), in order to properly explore the pesticides chemical space. To avoid data leakage, we explain only the model predictions on molecules from the test set.
@@ 3 Data and software availability
We release our code publicly at https://github.com/j-adamczyk/ApisTox_bee_toxicity_ML_prediction. The ApisTox dataset, along with MaxMin and time splits, is available on Zenodo [93] and GitHub [94]. We also include it in our GitHub repository with code for reproducing experimental results.
@@ 4 Results and discussion
The main results and quality metrics are reported in Table 1 (MaxMin split) and Table 2 (time split). For brevity, we include the top 5 fingerprints (with the highest MCC scores), with results for the rest in the B. The best metric value in each group is marked in bold (two values in case of ties).
@@
Molecular fingerprints are, overall, the most performant and robust solution. Interestingly, quite different fingerprints are the best for both splits. For MaxMin, the best fingerprint is Avalon [95], with 3 substructural fingerprints (Laggner [20], MACCS [21], PubChem [24]) also among the top 5. However, for time split, all best fingerprints are hashed, e.g. ECFP [23], Layered [92], RDKit [92]. This confirms that flexible, more data-driven hashed fingerprints are better suited for out-of-distribution generalization. In contrast, in-domain predictions can be made accurately by extracting common functional groups and other substructural patterns. Since a very large number of fingerprint algorithms are available, understanding such implications can be useful for practical applications.
The performance of GNNs trained from scratch is very underwhelming in both cases. All are unable to outperform baselines, even the simplest atom count in the MaxMin split. This is despite significant computational resources and extensive hyperparameter tuning, even for AttentiveFP, designed specifically for molecular property prediction. However, GNNs are known to require larger amounts of data to train effectively, and by its nature, agrochemistry requires low-data ML. Furthermore, it frequently involves salts, mixtures, and other molecules with disconnected components in their molecular graphs [14, 96, 97]. This is problematic for message-passing models, since they can obtain this information only in the last readout layer, limiting the effectiveness of node embedding learning from their neighborhoods.
@@ 4.1.3 Share of fragmented molecules
Fragmented molecules are defined as those in which two or more independent fragments are present in a single SMILES record. Examples include salts or mixtures, frequently found in agrochemical QSAR [98, 99]. Such compounds can be challenging for some methods, especially those based on graph neural networks (GNNs) or path-based fingerprints like Atom Pair. In such cases, the individual fragments are treated as separate graphs, which prevents the message passing between them. This can impact the performance of the models.
@@ 4.1.4 Unique functional groups
The functional groups defined in the Laggner fingerprint [20] were utilized. It is a collection of 307 fragments, also including those typically indicating toxicity or generally less common in typical drug-like collections.
@@ 4.1.5 Chemical diversity
Furthermore, all analyzed datasets appear to be relatively distant from each other (average Tanimoto similarity $\leq 0.56$), suggesting that they represent different areas of chemical space. However, it should be noted that the results obtained from this analysis show little correlation with the difficulty of their respective classification tasks and the quality of the resulting ML models [34].
@@ 4.2 Interpretability - counterfactual explanations
Overall, the most prevalent motifs influencing model predictions are connected to polarity, ionization states, charge distribution, and stability (e.g. ester bonds). In fact, these are some of the key factors that organic chemists and environmental scientists consider when designing molecules. The polarity and ionization states influence solubility and control the mobility of the compound in the environment, affecting aspects like runoff, leaching into groundwater, and bioavailability. More polar molecules tend to have lower bioaccumulation potential, as they interact less strongly with the lipid membranes of organisms [100]. Charge characteristics can dictate how a molecule adsorbs in soils or sediments, affecting its overall persistence and exposure risks [101]. Finally, incorporating chemically labile groups like ester bonds is a common strategy to design molecules that break down into less harmful components after they have performed their intended function. This is part of a “benign-by-design” approach to reduce long-term environmental impact [102].
@@ 4.3 Interpretability - uncertain molecules
In addition to the main dataset and its splits, ApisTox provides a smaller dataset of molecules rejected from the main dataset due to disagreements between experimental measurements in the ECOTOX database. For the main dataset, only molecules with perfectly agreeing $LD_{50}$ values were included, i.e., all under or over $11$ $\mu\text{g / bee}$ (US EPA guidelines threshold [103]). Others are provided as a file with uncertain molecules. Performing predictions on those compounds provides additional evidence, in addition to the literature and experiments, about their toxicity. This is also a very practical application of proposed models, as conflicting measurements can easily arise in ecotoxicology research [104, 105], e.g. due to variations in experimental procedures or outliers in results.
@@
Indoxacarb, also known commercially as Steward, Avaunt, Advion, and Arilon, is an oxadiazine insecticide, used in many household pesticides, and is also approved for use in the US and UK. However, the EU withdrew its approval for the use of Indoxacarb as a plant protection insecticide at the end of 2021, due to its environmental impact and high toxicity. In particular, “a high risk to bees was identified for the representative use in maize, sweet corn and lettuce for seed production” [106]. Thus, this is a rather nuanced situation with regulatory disagreements, and prediction of machine learning further corroborates the high toxicity of Indoxacarb toward honey bees.
Phosalone has values ranging from $9.9$ to $17.4$ in the data. The model predicts toxicity (positive class), which is in agreement with the EU decision from 2006 [107].